OpenStudy (anonymous):

Write each expression as a single logarithm. Log\/3 24 + Log\/3 2

OpenStudy (anonymous):



OpenStudy (anonymous):

\(\log_{3}{24}+\log_{3}{2}=\log_{3}{24 \cdot 2}=\log_{3}{48}\)

OpenStudy (anonymous):

Sorry I don't know why the two is so high. It's suppose to be even with the rest of it besides the three's

OpenStudy (anonymous):

Oh! Thank you! That's a lot easier than how It was explained to me. Thank you!

OpenStudy (anonymous):

your welcome! :)

OpenStudy (anonymous):

What do you do if it is like this 3 Logx + 5 Logy

OpenStudy (anonymous):

First, by the power rule, \(3\log{x}+5\log{y}=\log{x^3}+\log{y^5}\). Then proceed as before.

OpenStudy (anonymous):

Then you multiply 3 * 5?

OpenStudy (anonymous):

Nope. You must multiply \(x^3\) and \(y^5\), something like \(\log{x^3 y^5}\).

OpenStudy (anonymous):

You cannot multiply 3 and 5, since their bases, x and y, are not the same.

OpenStudy (anonymous):

So after you multiply them what do you do? Because wouldn't you still have x and y left

OpenStudy (anonymous):

That's the answer. x and y are not the same, so you cannot combine the two together. You can just put it side by side, like \(\log_{5}{x^4 y^7 z^9}\).

OpenStudy (anonymous):

So it would just be logx\[Logx ^{3}y ^{5} \] That's the answer?

OpenStudy (anonymous):

yes. :)

OpenStudy (anonymous):

Oh! Okay thanks again.